OpenStudy (anonymous):

2x/x+1= 3/x-2 + 1/ x^2-x-2

OpenStudy (anonymous):

now when doing this do I need to move things because of the = sign?

OpenStudy (pfenn1):

Am I reading this equation correctly? \[\frac {2x}{x+1}= \frac{3}{x-2} + \frac{1}{x^2-x-2}\]

OpenStudy (anonymous):

yes, for some reason it gives me an error when i use the equation thing

OpenStudy (pfenn1):

well, can you factor\[x^2-x-2\]

OpenStudy (anonymous):

so would it go down to just x^2 on the bottom?

OpenStudy (pfenn1):

no. Do you know what I mean by factor? what values of a and b would solve this equation?\[x^2-x-2=(x+a)(x+b)\] where we know that a + b = -1 and ab = -2

OpenStudy (pfenn1):

We are trying to find the LCD (lowest common denominator).

OpenStudy (anonymous):